Calculate the pH of a solution that is 0.253 M in nitrous acid (HNO2) and 1.687 M in potassium nitrite (KNO2). The Ka value of nitrous acid is 4.50 × 10-4.


1
Expert's answer
2021-09-09T02:02:01-0400

Ka = [H+][NO2-]/[HNO2]

4.5x10-4 = x(1.687)/.253 X = 3.0 x 10-3

pH = -log 3.0 x 10-3 =2.5